A 310 second-song is divided into three parts, the beginning, the middle, and the end. The beginning is one-sixth as long as the end. The middle part is one-half as long as the end. Find the length of each part
Answer by rfer(16322)

1/6x+1/2x+x=310
1/6x+3/6x+x=310
1 4/6x=310
10/6x=310
x=310/1*6/10
x=1860/10
x=186 sec
1/6x=31 sec
1/2x=93 sec